Find three consecutive numbers such that the sum of one-fourth the first and one-fifth the second is five less than one-seventh the third.
Answer by edjones(7569)

Let the numbers be x, x+1, x+2.
x/4 + (x+1)/5 = (x+2)/7 - 5
35x+28(x+1)=20(x+2)-700 Multiply each side by 140 (LCM).
35x+28x+28=20x+40-700
43x=-688
x=-16, -15, -14 the three numbers.
.
Ed